# Implicit geometric regularization in flow matching via density weighted Stein operators

### v1

@@ Abstract
Flow Matching (FM) has emerged as a powerful paradigm for continuous normalizing flows, yet standard FM implicitly performs an unweighted $L^{2}$ regression over the entire ambient space. In high dimensions, this leads to a fundamental inefficiency: the vast majority of the integration domain consists of low-density “void” regions where the target velocity fields are often chaotic or ill-defined. In this paper, we propose $\gamma$-Flow Matching ($\gamma$-FM), a density-weighted variant that aligns the regression geometry with the underlying probability flow. While density weighting is desirable, naive implementations would require evaluating the intractable target density. We circumvent this by introducing a Dynamic Density-Weighting strategy that estimates the target density directly from training particles. This approach allows us to dynamically downweight the regression loss in void regions without compromising the simulation-free nature of FM. Theoretically, we establish that $\gamma$-FM minimizes the transport cost on a statistical manifold endowed with the $\gamma$-Stein metric. Spectral analysis further suggests that this geometry induces an implicit Sobolev regularization, effectively damping high-frequency oscillations in void regions. Empirically, $\gamma$-FM significantly improves vector field smoothness and sampling efficiency on high-dimensional latent datasets, while demonstrating intrinsic robustness to outliers.
@@ 1 Introduction
The Manifold Hypothesis posits that high-dimensional real-world data concentrate near a low-dimensional manifold embedded in the ambient space (Fefferman2016Testing). While theoretical verification of this hypothesis remains a subject of active research (Pope2021Intrinsic), the remarkable success of deep generative models offers a constructive validation: if data were uniformly distributed in the high-dimensional void, efficient learning of the probability distribution would be computationally intractable (Bengio2013Representation). Thus, the capability to accurately model and sample from the data distribution is, in itself, a testament to the existence of such low-dimensional structures.
@@
Flow Matching (FM) has emerged as a powerful paradigm for capturing this distribution by unifying diffusion models and continuous normalizing flows (CNFs) (Lipman2023; albergo2023stochastic). It directly regresses a vector field that transports a simple base distribution to the data distribution. Unlike score-based diffusion models, FM avoids the need to estimate the score function or to solve reverse-time stochastic differential equations. Instead, it learns a deterministic ordinary differential equation (ODE) whose solution defines an invertible map between latent noise and data.
@@
Despite its conceptual elegance, FM faces a fundamental challenge in high-dimensional settings: the curse of dimensionality and volume imbalance. In high dimensions, the data manifold occupies a negligible fraction of the ambient space, and the majority of the integration domain consists of low-density “voids.” In standard FM, the training objective uniformly integrates the regression error over the entire probability path. Consequently, the model is forced to solve the regression problem even in these vast void regions, where probability paths are sparse or effectively irrelevant to the final generation quality. Forcing a neural network to fit target velocities in these “don’t-care” regions results in a rough vector field that wastes model capacity and causes numerical stiffness during ODE integration.
@@
In this work, we propose $\gamma$-Flow Matching ($\gamma$-FM) as a principled remedy for this issue. Our key idea is to reinterpret FM as a regression problem under a density-weighted geometry induced by the $\gamma$-divergence (Fujisawa2008). Instead of treating all spatial locations equally, we utilize the power density $p_{t}(x)^{\gamma}$ as an importance weight that naturally highlights the data manifold. This approach effectively “focuses” the learning process: the model prioritizes accurate vector field estimation where the data actually resides, while being allowed to remain smooth and simple in the empty ambient space. Latent flow matching has been independently explored by Lipman2023, who combine standard flow matching with pre-trained autoencoders and provide a Wasserstein-2 control for the resulting latent flows. Our work is complementary: we adopt a similar latent-flow setting, but instead of modifying the representation, we modify the regression geometry itself via the $\gamma$-weighted objective and analyse its effect through a $\gamma$-Stein and nonlinear Fokker–Planck viewpoint. While chen2024riemannian explicitly extend Flow Matching to Riemannian manifolds, our approach induces an implicit manifold geometry in the ambient space via density weighting, avoiding the need for explicit charts or geodesics.
@@
    Dynamic Density-Weighting: We formulate $\gamma$-FM as a weighted regression scheme that leverages the target density $p_{t}$. We propose a tractable, simulation-free implementation using particle-based density estimation.
@@
    Variance Reduction Analysis: We theoretically demonstrate that our weighting scheme minimizes the variance of the gradient estimator by suppressing high-noise signals in low-density regions.
@@
    Geometric Regularization: We show that $\gamma$-FM implicitly regularizes the vector field, reducing the Jacobian norm and enabling high-quality sampling with fewer function evaluations (NFE).
@@
    Robustness & Efficiency: Experiments on latent space CIFAR-10 confirm that $\gamma$-FM achieves a superior trade-off between generation quality and computational cost, while exhibiting strong robustness to outliers.
@@ 2.1 Conditional Flow Matching
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1]}\mathbb{E}_{x_{t}\sim p_{t}(x)}\left[\|v_{\theta}(x_{t},t)-u_{t}(x_{t})\|^{2}\right].$
@@
To solve this, Lipman2023 introduced Conditional Flow Matching (CFM), which uses a conditional probability path $p_{t}(x\mid x_{1})$ given the data endpoint $x_{1}$, and a corresponding conditional vector field $u_{t}(x\mid x_{1})$. Crucially, it has been shown that the gradients of the intractable objective (2.1) are identical to those of the conditional objective:
@@ 2.2 Motivation: Geometric Focusing via $\gamma$-Divergence
From the information-geometric viewpoint, it is natural to interpret flow matching through the geometry induced by divergences and the associated Riemannian metrics on statistical manifolds (Amari2000methods; ay2017information). Standard Flow Matching implicitly minimizes the discrepancy between the target and model vector fields in an unweighted $L^{2}$ sense. From the perspective of Optimal Transport, this objective corresponds to minimizing the standard kinetic energy of the flow, $\mathcal{E}(v)=\int\|v(x)\|^{2}p_{t}(x)\,\mathrm{d}x$, which is associated with the Wasserstein-2 geometry and the linear heat equation. Recent works have explicitly targeted Wasserstein-optimal paths via minibatch optimal transport couplings (tong2024improving), yet these approaches typically operate in the unweighted $L^{2}$ geometry. While statistically consistent, this "flat" geometry is inefficient in high dimensions because it assigns equal transport cost to the vast low-density "voids" as it does to the concentrated data manifold. To address this, we propose changing the underlying geometry of the regression itself, moving from the L2 regression (kinetic-energy) viewpoint to the robust $\gamma$-geometry, in analogy with the transition from the Fisher divergence to the $\gamma$-Fisher divergence; see barp2019minimum for the diffusion score-matching divergence.
@@
We define the $\gamma$-Flow Matching ($\gamma$-FM) objective as the minimization of the weighted kinetic energy. Instead of the standard energy, we align the regression with the density-weighted geometry induced by the $\gamma$-divergence (Fujisawa2008). We define the weighted loss as:
@@
This defines a Riemannian metric (specifically, the $\gamma$-weighted Fisher information metric) where the infinitesimal transport cost is scaled by the density power $p^{1+\gamma}$. In this geometry, distances in low-density regions are compressed to zero. Consequently, $\gamma$-FM does not simply "ignore" outliers; it solves the regression problem on a statistical manifold where the voids are geometrically insignificant. This modification naturally links the regression to the Porous Medium Equation (nonlinear diffusion) rather than the Heat Equation (linear diffusion), providing a theoretical guarantee for compact support preservation as discussed in Section 3, see Otto2001 for extensive discussion.
@@ 2.3 Tractability via Particle-Based Estimation
We circumvent this bottleneck by adopting a particle-based estimation strategy. Since the mini-batch samples $\mathcal{B}_{t}=\{x_{t}^{(i)}\}_{i=1}^{B}$ at any given time $t$ are drawn from $p_{t}$, their spatial distribution provides a direct Monte Carlo estimate of the density. We employ a robust kernel-based proxy for the density. Specifically, we define the weight for a sample $x_{t}\in\mathcal{B}_{t}$ based on the distance to its $k$-nearest neighbors:
@@ 3 Theoretical Analysis
In this section, we analyze the theoretical properties of $\gamma$-FM. We establish that our density-weighting scheme is not merely a heuristic, but acts as a variance-optimal estimator and enforces physical constraints consistent with nonlinear diffusion. To provide a concrete basis for the following analysis, we summarize the practical training procedure of $\gamma$-FM in Algorithm 1. This algorithm implements the simulation-free density estimation discussed in Section 2. For simplicity we use linear interpolants in experiments.
@@
8:  Sample time steps $t\sim\mathcal{U}[0,1]$
@@
10:  $x_{t}\leftarrow(1-t)x_{0}+tx_{1}$
@@
13:  if $\gamma>0$ then
14:   Compute pairwise distances matrix for batch $\{x_{t}\}$
@@
24:  $\mathcal{L}(\theta)\leftarrow\frac{1}{B}\sum_{i=1}^{B}w_{i}\|v_{\theta}(x_{t}^{(i)},t)-u_{t}^{(i)}\|^{2}$
@@ 3.2 Physical Consistency with Liouville Dynamics
Notice that this functional corresponds exactly to the self-divergence term in the definition of the $\gamma$-divergence (up to a sign). Just as minimizing $\gamma$-divergence statistically ignores outliers (Section 3), the gradient flow of its associated entropy $\mathcal{F}_{\gamma}$ physically restricts the spread of probability mass.
The Wasserstein gradient flow is defined by the continuity equation driving the density along the gradient of the variation (see, e.g., Jordan1998; AmbrosioGigliSavare2008).
@@
$\partial_{t}p(x,t)=\frac{\gamma+1}{\gamma}\nabla\cdot(p(x,t)\nabla p(x,t)^{\gamma})=\Delta(p(x,t)^{1+\gamma}),$ (8)
@@
The proof relies on the Comparison Principle for the Porous Medium Equation (vazquez2007porous). The principle states that if two solutions $u$ and $v$ satisfy $u(x,0)\leq v(x,0)$ everywhere, then $u(x,t)\leq v(x,t)$ for all $t>0$. We construct an explicit supersolution using the Barenblatt–Pattle solution $B(x,t;M)$, which represents the diffusion of a Dirac mass $M$. Let $B(x,\tau;M)$ be the Barenblatt profile centered at the origin with mass $M$ at a time shift $\tau>0$:
$B(x,\tau)=\tau^{-\alpha}\left(C(M)-\kappa\tau^{-2\beta}|x|^{2}\right)_{+}^{1/\gamma},$
@@
is, up to an affine rescaling, the negative of the Tsallis $q$-entropy with $q=1+\gamma$ (Tsallis1988). It is well known that maximizing the Tsallis entropy under mass and second-moment constraints,
@@
for suitable parameters $\beta>0$ and $\mu\in\mathbb{R}^{d}$. These $q$-Gaussians are precisely the self-similar Barenblatt profiles of the porous medium equation (8) for an appropriate correspondence between $q$ and the nonlinearity exponent $1+\gamma$; see, e.g., Malacarne2001; Takatsu2012. In particular, they exhibit compact support when $\gamma>0$, providing concrete examples of the finite-support behaviour described in Proposition 3.2.
@@ 3.3 Motivating Example: 1D Double-Well Potential
$\partial_{t}p_{t}=\nabla\cdot(p_{t}\nabla V)+D\nabla\cdot(p_{t}\nabla p_{t}^{\gamma}),$ (9)
@@
This analytical example provides a rigorous justification for the "Manifold Focusing" effect: our dynamic weighting $w_{\gamma}$ empirically emulates the finite-propagation physics of the PME, naturally enforcing compact support for the learned distribution.

### v2

@@ Abstract
Flow Matching (FM) has emerged as a powerful paradigm for continuous normalizing flows, yet standard FM implicitly performs an unweighted $L^{2}$ regression over the entire ambient space. In high dimensions, this leads to a fundamental inefficiency: the vast majority of the integration domain consists of low-density “void” regions where the target velocity fields are often chaotic or ill-defined. In this paper, we propose $\gamma$-Flow Matching ($\gamma$-FM), a density-weighted variant that aligns the regression geometry with the underlying probability flow. While density weighting is desirable, naive implementations would require evaluating the intractable target density. We circumvent this by introducing a Dynamic Density-Weighting strategy that estimates the target density directly from training particles. This approach allows us to dynamically downweight the regression loss in void regions without compromising the simulation-free nature of FM. Theoretically, we formulate an ideal $\gamma$-weighted regression geometry motivated by the $\gamma$-Stein metric, derive a variance-suppression bound for low-density regions, and use a weighted Dirichlet/spectral analysis to suggest a mechanism for smoother learned vector fields. Empirically, we evaluate $\gamma$-FM under a shared-time density-estimation protocol and compare it against both standard FM and an explicit Jacobian-regularized baseline using latent-space and image-space metrics.
@@ 1 Introduction
The Manifold Hypothesis posits that high-dimensional real-world data concentrate near a low-dimensional manifold embedded in the ambient space (Fefferman et al., 2016). While theoretical verification of this hypothesis remains a subject of active research (Pope et al., 2021), the remarkable success of deep generative models offers a constructive validation: if data were uniformly distributed in the high-dimensional void, efficient learning of the probability distribution would be computationally intractable (Bengio et al., 2013). Thus, the capability to accurately model and sample from the data distribution is, in itself, a testament to the existence of such low-dimensional structures.
@@
Flow Matching (FM) has emerged as a powerful paradigm for capturing this distribution by unifying diffusion models and continuous normalizing flows (CNFs) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023). It directly regresses a vector field that transports a simple base distribution to the data distribution. Unlike score-based diffusion models, FM avoids the need to estimate the score function or to solve reverse-time stochastic differential equations. Instead, it learns a deterministic ordinary differential equation (ODE) whose solution defines an invertible map between latent noise and data.
@@
Despite its conceptual elegance, FM faces a fundamental challenge in high-dimensional settings: the curse of dimensionality and volume imbalance. In high dimensions, the data manifold occupies a negligible fraction of the ambient space, and the majority of the integration domain consists of low-density “voids.” In standard FM, the training objective uniformly integrates the regression error over the entire probability path. Consequently, the model is forced to solve the regression problem even in these vast void regions, where probability paths are sparse or effectively irrelevant to the final generation quality. Forcing a neural network to fit target velocities in these “don’t-care” regions can result in a rough vector field and increased numerical stiffness during ODE integration.
@@
In this work, we propose $\gamma$-Flow Matching ($\gamma$-FM) as a principled remedy for this issue. Our key idea is to reinterpret FM as a regression problem under a density-weighted geometry induced by the $\gamma$-divergence (Fujisawa and Eguchi, 2008). Instead of treating all spatial locations equally, we utilize the power density $p_{t}(x)^{\gamma}$ as an importance weight that naturally highlights the data manifold. This approach effectively “focuses” the learning process: the model prioritizes accurate vector field estimation where the data actually resides, while being allowed to remain smooth and simple in the empty ambient space. Latent flow matching has been independently explored by Lipman et al. (2023), who combine standard flow matching with pre-trained autoencoders and provide a Wasserstein-2 control for the resulting latent flows. Our work is complementary: we adopt a similar latent-flow setting, but instead of modifying the representation, we modify the regression geometry itself via the $\gamma$-weighted objective and analyse its effect through a $\gamma$-Stein and nonlinear Fokker–Planck viewpoint. While Chen and Lipman (2024) explicitly extend Flow Matching to Riemannian manifolds, our approach induces an implicit manifold geometry in the ambient space via density weighting, avoiding the need for explicit charts or geodesics.
@@
    Shared-time dynamic density-weighting: We formulate $\gamma$-FM as a density-weighted regression scheme motivated by $p_{t}(x)^{\gamma}$ and implement it with a simulation-free particle surrogate computed from a shared-time minibatch, so that the $k$-NN statistics are aligned with a single marginal $p_{t}$.
@@
    Formal results versus mechanisms: We provide a variance-suppression bound showing how density weighting downweights high-variance updates from low-density regions, while clearly separating this formal statement from the continuum analogy and from the spectral mechanism used for intuition.
@@
    Geometric and spectral interpretation: We connect the idealized $\gamma$-weighted objective to $\gamma$-Stein/escort geometry and show, through a weighted Dirichlet viewpoint, how such geometry is naturally associated with smoother vector fields in populated regions.
@@
    Empirical protocol: We strengthen the latent-CIFAR evaluation with a matched-budget comparison against FM++Jac, and with decoded-image metrics including FID, Recall, and Coverage in addition to latent-space discrepancy, smoothness, and NFE.
@@ 2.1 Conditional Flow Matching
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1]}\mathbb{E}_{x_{t}\sim p_{t}(x)}\left[\|v_{\theta}(x_{t},t)-u_{t}(x_{t})\|^{2}\right].$ (1)
@@
To solve this, Lipman et al. (2023) introduced Conditional Flow Matching (CFM), which uses a conditional probability path $p_{t}(x\mid x_{1})$ given the data endpoint $x_{1}$, and a corresponding conditional vector field $u_{t}(x\mid x_{1})$. Crucially, it has been shown that the gradients of the intractable objective (1) are identical to those of the conditional objective:
@@ 2.2 Motivation: Geometric Focusing via $\gamma$-Divergence
From the information-geometric viewpoint, it is natural to interpret flow matching through the geometry induced by divergences and the associated Riemannian metrics on statistical manifolds (Amari and Nagaoka, 2000; Ay et al., 2017). Standard Flow Matching implicitly minimizes the discrepancy between the target and model vector fields in an unweighted $L^{2}$ sense. From the perspective of Optimal Transport, this objective corresponds to minimizing the standard kinetic energy of the flow, $\mathcal{E}(v)=\int\|v(x)\|^{2}p_{t}(x)\,\mathrm{d}x$, which is associated with the Wasserstein-2 geometry and the linear heat equation. Recent works have explicitly targeted Wasserstein-optimal paths via minibatch optimal transport couplings (Tong et al., 2024), yet these approaches typically operate in the unweighted $L^{2}$ geometry. While statistically consistent, this "flat" geometry is inefficient in high dimensions because it assigns equal transport cost to the vast low-density "voids" as it does to the concentrated data manifold. To address this, we propose changing the underlying geometry of the regression itself, moving from the L2 regression (kinetic-energy) viewpoint to the robust $\gamma$-geometry, in analogy with the transition from the Fisher divergence to the $\gamma$-Fisher divergence; see Barp et al. (2019) for the diffusion score-matching divergence.
@@
We define the $\gamma$-Flow Matching ($\gamma$-FM) objective as the minimization of the weighted kinetic energy. Instead of the standard energy, we align the regression with the density-weighted geometry induced by the $\gamma$-divergence (Fujisawa and Eguchi, 2008). We define the weighted loss as:
@@
This defines a Riemannian metric (specifically, the $\gamma$-weighted Fisher information metric) where the infinitesimal transport cost is scaled by the density power $p^{1+\gamma}$. In this geometry, distances in low-density regions are compressed to zero. Consequently, $\gamma$-FM does not simply "ignore" outliers; it solves the regression problem on a statistical manifold where the voids are geometrically insignificant. This modification naturally links the regression to the Porous Medium Equation (nonlinear diffusion) rather than the Heat Equation (linear diffusion), providing a theoretical guarantee for compact support preservation as discussed in Section 3, see Otto (2001) for extensive discussion.
@@ 2.3 Tractability via Particle-Based Estimation
We circumvent this bottleneck by adopting a particle-based estimation strategy. In the implementation, each minibatch uses a single shared time $t_{\mathrm{batch}}\sim\mathcal{U}[0,1]$, and we form $\mathcal{B}_{t_{\mathrm{batch}}}=\{x_{t_{\mathrm{batch}}}^{(i)}\}_{i=1}^{B}$. Thus the particle cloud used in the $k$-NN step is a Monte Carlo sample from the single marginal $p_{t_{\mathrm{batch}}}$, rather than a mixture over times, and its spatial distribution provides the intended local density proxy. We employ a robust kernel-based proxy for the density. Specifically, we define the weight for a sample $x_{t}\in\mathcal{B}_{t}$ based on the distance to its $k$-nearest neighbors:
@@ 3 Theoretical Analysis
In this section, we separate three layers of the argument: (i) formal statements that apply directly to the stated weighted objective, (ii) continuum analogies that clarify why low-density weighting suppresses motion in void regions, and (iii) spectral/Dirichlet mechanisms that help interpret the observed smoothing effect. To provide a concrete basis for the following analysis, we summarize the practical shared-time training procedure of $\gamma$-FM in Algorithm 1. This algorithm implements the simulation-free density estimation discussed in Section 2. For simplicity we use linear interpolants in experiments.
@@
8:  Sample a shared minibatch time $t_{\mathrm{batch}}\sim\mathcal{U}[0,1]$
@@
10:  $x_{t}\leftarrow(1-t_{\mathrm{batch}})x_{0}+t_{\mathrm{batch}}x_{1}$
@@
13:  Compute pairwise distances matrix within the shared-time batch $\{x_{t}\}$
14:  if $\gamma>0$ then
@@
24:  $\mathcal{L}(\theta)\leftarrow\frac{1}{B}\sum_{i=1}^{B}w_{i}\|v_{\theta}(x_{t}^{(i)},t_{\mathrm{batch}})-u_{t}^{(i)}\|^{2}$
@@ 3.2 Physical Consistency with Liouville Dynamics
Notice that this functional corresponds exactly to the self-divergence term in the definition of the $\gamma$-divergence (up to a sign). Just as minimizing $\gamma$-divergence statistically ignores outliers as in the variance-reduction argument above, the gradient flow of its associated entropy $\mathcal{F}_{\gamma}$ physically restricts the spread of probability mass.
The Wasserstein gradient flow is defined by the continuity equation driving the density along the gradient of the variation (see, e.g., Jordan et al. (1998); Ambrosio et al. (2008)).
@@
$\partial_{t}p(x,t)=\frac{\gamma+1}{\gamma}\nabla\cdot(p(x,t)^{\gamma}\,\nabla p(x,t))=\Delta(p(x,t)^{1+\gamma}),$ (9)
@@
The proof relies on the Comparison Principle for the Porous Medium Equation (Vázquez, 2007). The principle states that if two solutions $u$ and $v$ satisfy $u(x,0)\leq v(x,0)$ everywhere, then $u(x,t)\leq v(x,t)$ for all $t>0$. We construct an explicit supersolution using the Barenblatt–Pattle solution $B(x,t;M)$, which represents the diffusion of a Dirac mass $M$. Let $B(x,\tau;M)$ be the Barenblatt profile centered at the origin with mass $M$ at a time shift $\tau>0$:
$B(x,\tau;M)=\tau^{-\alpha}\left(C(M)-\kappa\tau^{-2\beta}|x|^{2}\right)_{+}^{1/\gamma},$
@@
is, up to an affine rescaling, the negative of the Tsallis $q$-entropy with $q=1+\gamma$ (Tsallis, 1988). It is well known that maximizing the Tsallis entropy under mass and second-moment constraints,
@@
for suitable parameters $\beta>0$ and $\mu\in\mathbb{R}^{d}$. These $q$-Gaussians are precisely the self-similar Barenblatt profiles of the porous medium equation (9) for an appropriate correspondence between $q$ and the nonlinearity exponent $1+\gamma$; see, e.g., Malacarne et al. (2001); Takatsu (2012). In particular, they exhibit compact support when $\gamma>0$, providing concrete examples of the finite-support behaviour described in Proposition 3.2.
@@ 3.3 Motivating Example: 1D Double-Well Potential
$\partial_{t}p_{t}=\nabla\cdot(p_{t}\nabla V)+D\nabla\cdot(p_{t}^{\gamma}\,\nabla p_{t}),$ (10)
@@
This analytical example should be read as a concrete illustration of the “Manifold Focusing” effect: our dynamic weighting $w_{\gamma}$ mimics the finite-propagation intuition of the PME and helps explain why the learned flow tends to avoid low-density voids, rather than as a derivation of the exact neural training dynamics.